OpenStudy (anonymous):

Solve the equation x2 − 100 = 0.

Parth (parthkohli):

Difference of squares. I see.

Parth (parthkohli):

\(a^2 - b^2 = (a + b)(a - b)\)

OpenStudy (anonymous):

is it 10 , -10 ?

Parth (parthkohli):

\( \color{Red}{\Rightarrow a^2 - b^2 = (a + b)( a- b) \Longrightarrow \color{Black}{x^2 - 10^2 = .....} }\)

Parth (parthkohli):

Yes, it's (10,-10)